OpenStudy (anonymous):

7-5x(under radical)=8

OpenStudy (mimi_x3):

\(\sqrt{7-5x}=8\)?

OpenStudy (anonymous):

yes

OpenStudy (mimi_x3):

\(7-5x=8^{2}\) are you able to do it now?

OpenStudy (mimi_x3):

\(7-5x=64\) => \(-5x= 54\) => \(x=...\)

OpenStudy (anonymous):

Just a quick tip, @Mimi_x3 [you'e doing a great job] but when you see these questions with solving equations with radicals that you have to perform the opposite operation.

OpenStudy (mimi_x3):

I did perform the opposite operation..

OpenStudy (anonymous):

so what is the correct way to do this problem?

OpenStudy (mimi_x3):

I just showed you!

OpenStudy (anonymous):

oh it didn't load. sorry